At 20°C, the carbon resistor in an electric circuit connected to a 4.0 V battery has a resistance of 221 . What is the current in the circuit when the temperature ofthe carbon rises to 79°C?
Answer #1
Temperature coefficient of resistance for carbon = 0.004041 = alpha
R = 221 Ohms
R' = R*(1+alpha*(79-20)) = (221)*(1+(0.004041*(79-20))) = 273.6906 Ohms
I = V/R' = 4/273.6906 = 0.0146 A
answered by: Ann james
Answer #2
4.0v battery
source: Electricity
answered by: Amjad kawah